OpenStudy (anonymous):

how to slove f(3)= 3h + 1? giving out testimonals , medals and fans .. FOR HELP ONLY

OpenStudy (akashdeepdeb):

f(3) = 3h+1 f(a) = 3(a)+1 f(b) = 3(b)+1 . . . . . f(100) = 3(100)+1 Try now. :)

OpenStudy (akashdeepdeb):

Any input in value of function will give you the answer. So f(3) = 3(3)+1 Are you understanding? :)

OpenStudy (anonymous):

LOL THANK .. soo it wouldd bee f(3)=10 ?

OpenStudy (akashdeepdeb):

:) yes.